OpenStudy (anonymous):

The velocity v of the flow of blood at a distance r form the central axis of an artery of radius R is v = k(R^2 - r^2) Where k is the constant of proportionality. Find the average rate of flow of blood along a radius of the artery. (Use 0 and R as the limits of integration.)

OpenStudy (anonymous):

∫ (0, R)

OpenStudy (anonymous):

Would I use the velocity formula I am confused because there is not t or time involved to use this function.

OpenStudy (dumbcow):

\[avg = \frac{k}{R} \int\limits_0 ^R R^{2} - r^{2} dr\]

OpenStudy (anonymous):

Could you explain this to me please?

OpenStudy (anonymous):

This would involve a constant speed k?

OpenStudy (dumbcow):

velocity here is given in terms of "r" distance from center of artery in general to find avg value over interval is: \[\frac{1}{b-a} \int\limits_a^b f(x) dx\]

OpenStudy (anonymous):

Okay thanks I just found this in my book I was unsure. Definition of the average value of a function on an interval. If f is integrable on the closed interval[a, b], then the average value of f on the interval is

OpenStudy (dumbcow):

yep, then make sure you treat "R" as a constant